OpenStudy (anonymous):

A bicycle race has 9 contestants. How many ways can the contestants finish first, second, and third in the race? A.9 B.504 C.3024 D.362,880

OpenStudy (anonymous):

9!/(9-3)! = 9!/6! = 9*8*7=504 B. 504

OpenStudy (anonymous):

thanks so much!